OpenStudy (anonymous):

How do you factor: (x^3) + (2x^2) -4x -8 ?

OpenStudy (anonymous):

look for x+2 in both pairs

OpenStudy (anonymous):

(x^3 + 2x^2) + (-4x - 8) = x^2 (x + 2) + -4 (x+2) = (x^2 - 4) (x + 2) = (x + 2) (x - 2) (x + 2)